# Explaining Humour Style Classifications: An XAI Approach to Understanding Computational Humour Analysis

### v1

@@
Explaining Humour Style Classifications: An XAI Approach to Understanding Computational Humour Analysis
Mary Ogbuka Kenneth m.kenneth22@imperial.ac.uk Algorithmic Human Development group, Department of Computing, Imperial College London, UK Foaad Khosmood foaad@calpoly.edu Computer Engineering Department, California Polytechnic State University, USA Abbas Edalat a.edalat@imperial.ac.uk Algorithmic Human Development group, Department of Computing, Imperial College London, UK
Summarium
@@
Explainable AI, Computational Humour, Humour Style Classification, Natural Language Processing, Machine Learning Interpretability, Digital Humanities
@@ 1 Introduction
Explainable AI (XAI) has emerged as a solution to address the opacity of traditional ML models by providing insights into their decision-making processes Lyu et al. (2024). Through techniques such as Local Interpretable Model-Agnostic Explanations (LIME) Ribeiro et al. (2016) and Shapley Additive Explanations (SHAP) Lundberg and Lee (2017), XAI enables researchers to identify key features driving predictions. This interpretability is especially crucial for recognising humour styles, as the interaction of linguistic patterns, emotional tones, and semantic nuances necessitates a thorough comprehension. For instance, identifying features such as sarcasm, sentiment contrasts, or emotions can provide more detailed information about why a certain text is categorised as affiliative or self-deprecating.
@@ 2.2 Interpretable Models for Humour Analysis
Chen et al. (2024) introduced TalkFunny, a Chinese humour response dataset with chain-of-humourannotations and mind maps explaining response generation. Their PLM-based framework, integrating interpretability tools, showed improved performance. While focused on humour generation, their approach supports the value of explainable components, aligning with our goals for interpretability in humour AI.
@@ 2.3 Explainable Approaches to Style Classification
While significant progress has been made in humour classification and XAI techniques, two significant gaps persist: most interpretable approaches focus on binary rather than style-specific classification, and current style-specific approaches lack comprehensive explanatory frameworks for linguistic and emotional feature interaction. Our framework addresses these gaps through a unified approach to humour style classification explanation, supporting both theoretical understanding and practical applications.
Table 1 summarises the key works discussed in this section, highlighting the authors’ specific tasks, datasets used, features extracted, classification approaches, XAI methods employed, and their achieved results. This overview demonstrates both the progress made in computational humour analysis and explainability techniques while highlighting the gaps our work addresses.
Tabula 1: Summary of Related Works
Pérez-Landa et al. (2021) Xenophobia
Experts Xenophobia
Database (EXD) (10,057 tweets)
Pitropakis Xenophobia Database
(PXD)(5814 tweets)
TF-IDF, Bag of words (BOW),
Word 2 vector (W2V),
Keywords, Sentiment,
Emotion, and Syntactic
C45, KNN, Rusboost,
UnderBagging,
Feature engineering
(keywords, sentiment,
emotions and syntactic
F1: 76.8% (EXD)
AUC: 0.864 (EXD)
F1: 73.4% (PXD)
AUC: 0.794 (PXD)
Chowdhury et al. (2021) Sentiment Twitter API (2-years period)
W2V, Sentiment Specific Word
Embedding (SSWE), GloVe,
FastText embeddings
Bi-directional LSTM LIME
ACC: 72%
F1: 72%
Ahmed et al. (2022) Sentiment Patient-authored text (15,044)
Fuzzy logic rules, GloVe,
Cosine Similarity, Keywords
Bi-directional LSTM
Attention with fuzzy
classification
Fuzzy logic rules F1: 89%
Ortega-Bueno et al. (2022) Irony, Satire
IroSvA’19 shared task (3000)
Barbieri’15-es (10,000)
Salas’17-mx (10,888)
HAHA’19 shared task (6000)
BERT embedding,
Multilingual Universal
Sentence encoding,
semantics, affective, incongruity,
stylistic and structural
Multiview Attention LSTM
Feature engineering
(semantics, affective,
incongruity, stylistic
and structural)
F1: 70.4% (IroSva’19)
F1: 95.7% (Barbieri’15))
F1: 96% (Salas’17)
F1: 80.6% (HAHA’19)
Zhang et al. (2017) Humour 16000 One-Liner (32,002)
Contextual knowledge,
Subjectivity, affective polarity,
ambiguity, incongruity
Feature engineering
(Contextual knowledge,
Subjectivity, polarity,
ambiguity, incongruity)
F1: 85%
Mann and Mikulandric (2024) Humour, Pun CLEF 2024 JOKER Tasks TF-IDF
Logistic regression,
Naive Bayes (NB),
SVM, MarianMTModel
F1: 83% (Pun)
F1: 61% (Humour)
De Marez et al. (2024) Humour SemEval 2021 Task 7
Polarity, emotions, offense,
subjectivity, hate, stance,
ambiguity, adult language
GA2M
Feature engineering
(Polarity, emotions,
offense, subjectivity,
hate, stance, ambiguity,
adult language)
F1: 85%
Mahajan and Zaveri (2024) Humour Yelp reviews Emotive, incongruity, intensity
Stacking-based ensemble,
NB, SVM, MLP,
majority-vote ensemble
Feature engineering
(Berger’s typology-based
incongruity features)
F1: 72.57%
Chen et al. (2024) Humour Chain-of-humour dataset word embeddings
Pre-trained language
models (PLMs)
Chain-of-thought F1: 87%
Abulaish and Kamal (2018) Sarcasm Twitter API (7 weeks tweet)
Self-referencing, hyperbolic,
part-of-speech
Decision tree,
Naive Bayes,
Bagging
F1: 94%
Kamal and Abulaish (2020) Humour
Pun of the Day (4,826),
1600 One-Liner (32.002),
Twitter API (20,000)
Self-deprecating pattern,
exaggeration, word embedding,
ambiguity, interpersonal effect,
phonetic style.
Random forest
F1: 62% - 87%
across datasets
Kenneth Ogbuka et al. (2024) Humour Styles Humour styles dataset Sentence embedding
NB, Randon forest,
XGBoost, DistilBERT
F1: 78.6%
@@ 3.1 Dataset and Classification Model
  • Self-enhancing: 1,181 unique words, lengths ranging from 5 to 35 words
  • Self-deprecating: 1,365 unique words, lengths ranging from 4 to 148 words
  • Affiliative: 1,376 unique words, lengths ranging from 5 to 229 words
  • Aggressive: 1,449 unique words, lengths ranging from 4 to 112 words
  • Neutral: 1,856 unique words, lengths ranging from 4 to 34 words
@@
Descriptio 1: Examples for Each Category in the Dataset
@@ 3.3.1 Linguistic Patterns
  • Rhyme: The repetition of similar sounds at the ends of words. Examples include sight and flight, sad and mad, cat and hat. In our analysis, we used the “pronouncinglibrary with the CMU Pronouncing Dictionary to detect rhymes. Although the dictionary effectively covers standard English pronunciations, it may not capture slang or neologisms present in the dataset. For instance, out of the 4,506 unique words in the dataset, 90.75% (4,089 words) were found in the CMU Pronouncing Dictionary. Unmatched examples include: ‘launchalot’, ‘aronofskys’, ‘500000’, ‘ahappy’, ‘admited’, ‘kanye’, ‘houseplant’, ‘idk’, ‘behaviours’
  • Alliteration: The repetition of initial consonant sounds in closely placed words, such as in Crazy cats create chaosAAaaOOoo|<«>»""
. Implemented using NLTK’s phonetic detection and spaCy’s tokenisation, the algorithm identifies words sharing initial phonemes via the CMU Pronouncing Dictionary. However, this approach has limitations:
  • It considers only the first phoneme, occasionally grouping words incorrectly (e.g., "when" and "one").
@@
: This category analyses deliberate manipulations of word meanings to create humour.
  • Puns: Words with similar pronunciations but different meanings, identified through phonetic similarity and semantic difference analysis. For instance, "A bicycle can’t stand on its own because it is two-tired." Here, the bolded word is a homophonic pun, replacing too-tired with two-tired. Our implementation detects homophonic puns using phonetic matching (using ‘pronouncing’) and meaning differentiation (using WordNet). Current limitations include:
@@
: Structural elements focus on how linguistic arrangements contribute to humour.
@@ 3.3.2 Affective Patterns
: Sarcasm detection was performed using a RoBERTa-based model trained for sarcasm identification, achieving 60.7% accuracy. The model provides binary classification (sarcastic/non-sarcastic) and probability scores, capturing subtle verbal irony often found in specific humour styles.
@@
: Sentiment analysis employed a multi-layered approach using RoBERTa (93.2% average accuracy) (Hartmann et al., 2023) and TextBlob models to assess:
@@
: Emotional content was analysed using a DistilBERT-based model that classifies text into six discrete emotion categories (joy, anger, sadness, fear, love, and surprise) with an accuracy of 93.8% and F1-score of 93.79%. This fine-grained emotional analysis reveals distinctive affective patterns across different humour styles.
@@ 3.3.3 Contrast Patterns
: We examined opposing emotional valences at two distinct levels:
@@
: We investigated semantic relationships and conflicts through:
@@ 4 Results and Discussion
This section presents analysis of the ALI+XGBoost model’s performance in humour style classification, followed by detailed examination of linguistic mechanisms and error patterns. We first evaluate the model’s classification metrics and confidence scores across different humour styles, then analyse the distinctive linguistic features characterising each style, and finally examine misclassification patterns to understand the model’s limitations.
@@ 4.1 Model Performance Analysis
Confidence scores also varied significantly across style(p < 0.0001): Self-enhancing humour had the highest average confidence (0.889), followed by neutral (0.887), self-deprecating (0.811), aggressive (0.780), and affiliative humour (0.748).
@@ 4.2.1 Syllable Complexity
Neutral content exhibited significantly higher syllabic complexity (mean: 1.355, SD: 0.245), suggesting more formal or sophisticated language use. In contrast, aggressive humour demonstrated the lowest complexity (mean: 1.148, SD: 0.118), indicating a preference for simpler, more direct language. Other humour styles showed intermediate complexity levels, with self-enhancing (mean: 1.206) and self-deprecating (mean: 1.192) humour showing similar patterns.
@@ 4.2.2 Semantic and Structural Elements
Affiliative humour exhibited the highest frequency of semantic conflicts (mean: 29.755, SD: 114.755), significantly surpassing other styles. This suggests frequent use of wordplay and unexpected combinations. Self-deprecating humour followed with a mean of 18.717 (SD: 47.830). In contrast, neutral and self-enhancing content demonstrated markedly lower frequencies of semantic conflicts (mean: 7.025, SD: 7.347, and mean: 7.557, SD: 7.274, respectively). Homonym usage patterns revealed significant differences between humorous and neutral content. Social forms of humour—affiliative and aggressive—showed the highest homonym frequencies (mean = 6.796, SD = 4.743 and mean = 6.667, SD = 3.772, respectively), while neutral content demonstrated significantly lower usage (mean = 2.662, SD = 1.987). This indicates that wordplay involving multiple word meanings is a common feature across humour styles.
Tabula 2: Complexity Statistics
@@ 4.2.3 Rhetorical Devices
Exaggeration patterns varied significantly across styles. Self-deprecating humour exhibited the highest frequency of exaggeration (mean = 1.848, SD = 1.738), followed by aggressive humour (mean = 1.456, SD = 1.593). Neutral and self-enhancing content showed significantly lower exaggeration frequencies (mean = 0.738, SD = 0.853 and mean = 0.885, SD = 0.950, respectively). These findings suggest that humour styles employ distinctive combinations of linguistic mechanisms. Social forms of humour (affiliative and aggressive) predominantly utilise wordplay devices, while self-directed styles demonstrate varied complexity patterns. Neutral content maintains higher linguistic complexity but shows reduced usage of rhetorical devices.
@@ 4.2.4 Mechanism Correlations and Interactions
Analysis of correlations between linguistic mechanisms revealed complex interaction patterns in humour construction (Table 3). We observed a hierarchy of associations ranging from strong correlations to weak or negligible relationships.
Primary Correlations
The strongest association was identified between semantic conflicts and rhyme (r = 0.95), indicating substantial co-occurrence of these mechanisms in humorous content. This relationship suggests that rhyming elements may systematically contribute to the creation of semantic conflicts, potentially enhancing humorous effects through structural-semantic interactions.
Wordplay Networks
Homonyms demonstrated significant correlations with multiple linguistic features, forming a network of moderate-strength associations. These included correlations with semantic conflicts (r = 0.66), rhyme (r = 0.61), and exaggeration (r = 0.58). The strength and consistency of these correlations suggest that multiple-meaning words may function as central anchoring elements in humour construction, facilitating the integration of various linguistic devices.
Tabula 3: Mechanism Correlations
pun_count rhyme_count has_alliteration self_reference_count exaggeration_count semantic_conflict_count homonym_count
pun_count 1 0.29 0.06 0.12 0.24 0.28
rhyme_count 0.29 1 0.05 0.15 0.29 0.95 0.61
has_alliteration 0.05 0.05 1 0.13 0.17 0.05 0.21
self_reference_count 0.06 0.15 0.13 1 0.20 0.15 0.28
exaggeration_count 0.12 0.29 0.17 0.20 1 0.37 0.58
semantic_conflict_count 0.24 0.95 0.05 0.15 0.37 1 0.66
homonym_count 0.28 0.61 0.21 0.28 0.58 0.66 1
@@
Several mechanisms demonstrated operational independence. Self-referential content exhibited consistently weak correlations with other features $(0.06\leq r\leq 0.28)$, suggesting it functions largely autonomously in humour construction. Similarly, alliteration showed minimal correlation with other mechanisms (r < 0.21 across all comparisons), indicating its role as an independent supplementary device rather than a core component of humour construction.
Pun-Related Patterns
Pun usage demonstrated selective associations, showing moderate correlations with sound-based features including rhyme (r = 0.29) and homonyms (r = 0.28). However, correlations with other mechanisms were notably weaker (r < 0.20). This pattern suggests that while puns integrate sound patterns and multiple meanings, they operate through relatively distinct linguistic pathways in humour generation. These correlation patterns reveal a hierarchical structure in humour construction, where certain mechanisms (particularly rhyme and semantic conflicts) demonstrate strong interdependence, while others (such as alliteration and self-reference) function more independently. This finding suggests that humour may be constructed through multiple parallel linguistic pathways rather than a single unified mechanism.

### v2

@@
Abstract
@@
Explainable AI; Computational Humour; Humour Style Classification; Natural Language Processing; Machine Learning Interpretability; Digital Humanities
@@ 1 Introduction
Explainable AI (XAI) has emerged as a solution to address the opacity of traditional ML models by providing insights into their decision-making processes (Lyu et al., 2024). Through techniques such as Local Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) and Shapley Additive Explanations (SHAP) (Lundberg and Lee, 2017), XAI enables researchers to identify key features driving predictions. This interpretability is especially crucial for recognising humour styles, as the interaction of linguistic patterns, emotional tones, and semantic nuances necessitates a thorough comprehension. For instance, identifying features such as sarcasm, sentiment contrasts, or emotions can provide more detailed information about why a certain text is categorised as affiliative or self-deprecating.
@@ 2.2 Interpretable Models for Humour Analysis
Chen et al. (2024) introduced TalkFunny, a Chinese humour response dataset with ”chain-of-humour” annotations and mind maps explaining response generation. Their PLM-based framework, integrating interpretability tools, showed improved performance. While focused on humour generation, their approach supports the value of explainable components, aligning with our goals for interpretability in humour AI.
@@ 2.3 Explainable Approaches to Style Classification
While significant progress has been made in humour classification and XAI techniques, two significant gaps persist: most interpretable approaches focus on binary rather than style-specific classification, and current style-specific approaches lack comprehensive explanatory frameworks for linguistic and emotional feature interaction. Our framework addresses these gaps through a unified approach to humour style classification explanation, supporting both theoretical understanding and practical applications. A comprehensive summary of related works and their contributions is presented in Table 5 (Appendix A).
@@ 3.1 Dataset and Classification Model
  • Self-enhancing: 1,181 unique words, with lengths ranging from 5 to 35 words, median of 13 and mode of 10 words
  • Self-deprecating: 1,365 unique words, with lengths ranging from 4 to 148 words, median of 15 and mode of 12 words
  • Affiliative: 1,376 unique words, with lengths ranging from 5 to 229 words, median of 15 and mode of 13 words
  • Aggressive: 1,449 unique words, with lengths ranging from 4 to 112 words, median of 15 and mode of 12 words
  • Neutral: 1,856 unique words, with lengths ranging from 4 to 34 words, median of 10 and mode of 9 words
While the length distributions show positive skewness with some long examples (particularly in affiliative humour), the central tendencies are quite similar across humour styles, with medians ranging from 10-15 words and modes from 9-13 words. The neutral category shows the most compact distribution (4-34 words), while humorous categories occasionally include longer examples. However, since the majority of instances cluster around similar lengths, these outliers are unlikely to impact classification performance significantly.
@@
Figure 1: Examples for Each Category in the Dataset
@@ 3.3.1 Linguistic Patterns
  • Rhyme: The repetition of similar sounds at the ends of words. Examples include sight and flight, sad and mad, cat and hat. In our analysis, we used the “pronouncing” library with the CMU Pronouncing Dictionary to detect rhymes. Although the dictionary effectively covers standard English pronunciations, it may not capture slang or neologisms present in the dataset. For instance, out of the 4,506 unique words in the dataset, 90.75% (4,089 words) were found in the CMU Pronouncing Dictionary. Unmatched examples include: ‘launchalot’, ‘aronofskys’, ‘500000’, ‘ahappy’, ‘admited’, ‘kanye’, ‘houseplant’, ‘idk’, ‘behaviours’
  • Alliteration: The repetition of initial consonant sounds in closely placed words, such as in ”Crazy cats create chaos”. Implemented using NLTK’s phonetic detection and spaCy’s tokenisation, the algorithm identifies words sharing initial phonemes via the CMU Pronouncing Dictionary. However, this approach has limitations:
    • It considers only the first phoneme, occasionally grouping words incorrectly (e.g., ”when” and ”one”).
@@
This category analyses deliberate manipulations of word meanings to create humour.
  • Puns: Words with similar pronunciations but different meanings, identified through phonetic similarity and semantic difference analysis. For instance, ”A bicycle can’t stand on its own because it is two-tired.” Here, the bolded word is a homophonic pun, replacing too-tired with two-tired. Our implementation detects homophonic puns using phonetic matching (using ‘pronouncing’) and meaning differentiation (using WordNet). Current limitations include:
@@
Structural elements focus on how linguistic arrangements contribute to humour.
@@ 3.3.2 Affective Patterns
Sarcasm detection was performed using a RoBERTa-based model trained for sarcasm identification, achieving 60.7% accuracy. The model provides binary classification (sarcastic/non-sarcastic) and probability scores, capturing subtle verbal irony often found in specific humour styles.
@@
Sentiment analysis employed a multi-layered approach using RoBERTa (93.2% average accuracy) (Hartmann et al., 2023) and TextBlob models to assess:
@@
Emotional content was analysed using a DistilBERT-based model that classifies text into six discrete emotion categories (joy, anger, sadness, fear, love, and surprise) with an accuracy of 93.8% and F1-score of 93.79%. This fine-grained emotional analysis reveals distinctive affective patterns across different humour styles.
@@ 3.3.3 Contrast Patterns
We examined opposing emotional valences at two distinct levels:
@@
We investigated semantic relationships and conflicts through:
@@ 4 Results and Discussion
This section presents an analysis of the ALI+XGBoost model’s performance in humour style classification, followed by a detailed examination of linguistic mechanisms and error patterns. We first evaluate the model’s classification metrics and confidence scores across different humour styles, then analyse the distinctive linguistic features characterising each style, and finally examine misclassification patterns to understand the model’s limitations. The complete LIME visualisations, linguistic analysis, and affective analysis results for all 293 examined instances are available in our public repository  [^2] XAI humour styles : https://github.com/MaryKenneth/XAI_humour_styles/tree/main.
@@ 4.1 Model Performance Analysis
Confidence scores also varied significantly across styles (p <0.0001): Self-enhancing humour had the highest average confidence (0.889), followed by neutral (0.887), self-deprecating (0.811), aggressive (0.780), and affiliative humour (0.748).
@@ 4.2.1 Syllable Complexity
Neutral content exhibited significantly higher syllabic complexity (mean: 1.355, SD: 0.245), suggesting more formal or complex language use. In contrast, aggressive humour demonstrated the lowest complexity (mean: 1.148, SD: 0.118), indicating a preference for simpler, more direct language. Other humour styles showed intermediate complexity levels, with self-enhancing (mean: 1.206) and self-deprecating (mean: 1.192) humour showing similar patterns.
@@ 4.2.2 Semantic and Structural Elements
Affiliative humour exhibited the highest frequency of semantic conflicts (mean: 29.755, SD: 114.755), significantly surpassing other styles. This suggests frequent use of wordplay and unexpected combinations. Self-deprecating humour followed with a mean of 18.717 (SD: 47.830). In contrast, neutral and self-enhancing content demonstrated markedly lower frequencies of semantic conflicts (mean: 7.025, SD: 7.347, and mean: 7.557, SD: 7.274, respectively).
Homonym usage patterns revealed significant differences between humorous and neutral content. Social forms of humour—affiliative and aggressive—showed the highest homonym frequencies (mean = 6.796, SD = 4.743 and mean = 6.667, SD = 3.772, respectively), while neutral content demonstrated significantly lower usage (mean = 2.662, SD = 1.987). This indicates that wordplay involving multiple word meanings is a common feature across humour styles.
@@ 4.2.3 Rhetorical Devices
Exaggeration patterns varied significantly across styles. Self-deprecating humour exhibited the highest frequency of exaggeration (mean = 1.848, SD = 1.738), followed by aggressive humour (mean = 1.456, SD = 1.593). Neutral and self-enhancing content showed significantly lower exaggeration frequencies (mean = 0.738, SD = 0.853 and mean = 0.885, SD = 0.950, respectively).
These findings suggest that humour styles employ distinctive combinations of linguistic mechanisms. Social forms of humour (affiliative and aggressive) predominantly utilise wordplay devices, while self-directed styles demonstrate varied complexity patterns. Neutral content maintains higher linguistic complexity but shows reduced usage of rhetorical devices.
@@ 4.2.4 Mechanism Correlations and Interactions
Analysis of correlations between linguistic mechanisms revealed complex interaction patterns in humour construction. We employed both Pearson and Spearman correlation analyses as shown in Figure 2 to understand these relationships more thoroughly, as Spearman correlations are more robust to outliers and can detect non-linear relationships.
Correlation Disparity and Outlier Effects
The most striking finding emerged in the disparity between Pearson and Spearman correlations, particularly for semantic conflict relationships. While Pearson correlation suggested a strong linear relationship between semantic conflicts and rhyme (r = 0.95), the much lower Spearman correlation (r = 0.32) revealed this relationship to be heavily influenced by outliers. This disparity manifests across different scaling levels:
  • High-density outliers: 794 conflicts/181 rhymes; 294 conflicts/44 rhymes; 231 conflicts/27 rhymes
  • Mid-range instances: 48 conflicts/7 rhymes; 30 conflicts/5 rhymes
  • Typical cases: Majority fall below 30 conflicts; Variable rhyme counts (0-5)
This distribution pattern is illustrated by two contrasting joke examples:
A typical case example: “Reminder that Winnie the Pooh wore a crop top with no pants and ate his fave food and loved himself. So you can too.”
  • 4 rhyming pairs: (“Pooh/you”, “Pooh/too”, “crop/top”, “you/too”)
  • 5 semantic conflict: memory aid vs physical action (“reminder/wore”), memory aid vs emotional satisfaction (“reminder/loved”), clothing vs eating (“wore/food”), physical vs emotional satisfaction (“food/loved”), and the phrase “So you can too” suggests an invitation, but applying it to “food” and “loved himself” creates ambiguity (“food/too”).
High rhyme example: “This is your mother. I just texted you but I don’t know how to make the facey-things so… happy face at the end.”
  • 4 distinct rhyming pairs (“you/do”, “you/to”, “do/to”, “know/so”)
  • No significant semantic conflicts
Figure 2: Mechanism Pearson and Spearman Correlations
Stable Correlations and Robust Relationships
While many correlations weakened under Spearman analysis, some relationships demonstrated stability across both methods. Most importantly, the relationship between rhyme and homonym remained consistent (Pearson: 0.61, Spearman: 0.62), suggesting a robust underlying connection independent of outliers. Homonyms also showed stable correlations with exaggeration (Pearson: 0.58, Spearman: 0.44) and semantic conflicts (Pearson: 0.66, Spearman: 0.30).
Consider this example that demonstrates stable mechanism interactions: “I forgot you can’t make depression jokes outside of twitter Imao my coworker was like ’you ready for this year to be over?’ I was like I’m ready for this life to be over’ he was like bro what”. This joke exhibits stable correlations between multiple features:
  • 7 rhyming pairs such as “I/my”, “you/to”, “be/he”
  • 6 homonyms like “I”: [“eye”, “aye”, “ai”]; ‘you”: [“yew”, “ewe”]
  • 10 semantic conflicts supporting wordplay such as ‘‘forgot/depression”, “forgot/year”, “forgot/life”, “make/year”
  • 3 self-references integrated with sound patterns-“I/my/I”
  • 12 alliteration instances like “F”: [“forgot”, “for”]; “Y”: [“you”, “year”]; “M”: [“make”, “my”]; “T”: [“Twitter”, ”to”]
@@
Several mechanisms demonstrated operational independence. Self-referential content maintained weak to moderate correlations with other features –Pearson range:$(0.06\leq r\leq 0.28)$ and similar stability range for Spearman, suggesting it functions largely autonomously in humour construction. Similarly, alliteration showed minimal correlation with other mechanisms (Pearson correlations <0.21 and Spearman correlations <0.27 across all comparisons), indicating its role as an independent supplementary device rather than a core component of humour construction.
Example joke: “My manager asked if I take constructive criticism and I said yes while wiping away my teary eyes.”
In the above joke, self-reference is present in four instances: “my/manager”, “I/take”, “I/said”, and “my/teary”. These create humour through a personal narrative, establishing a relatable and emotional context without requiring support from other mechanisms.
Alliteration appears in six groups, such as the “M” sounds in [“my,” “manager,” “my”] and the “T” sounds in [“take,” “teary”]. While alliteration enhances the rhythmic flow of the joke, it functions as a supplementary device, adding aesthetic value without depending on other features for its effect.
The semantic structure of the joke contains 25 semantic conflicts such as “manager/criticism”, “manager/teary”, and “manager/yes”, which operate independently of sound patterns. Looking at these three example conflicts: 1) the manager’s formal criticism meeting an emotional response - demonstrating the professional-personal divide that operates independently of phonetic features; 2) the stark contrast between professional authority and teary vulnerability - showcasing how semantic conflicts can generate humour without relying on sound patterns; and 3) the ironic juxtaposition of verbal affirmation against visible distress - illustrating how meaning-based conflicts drive the humour independently of linguistic devices like rhyme or wordplay.
This example illustrates how different humour mechanisms (self-reference, alliteration, and semantic conflict) can function independently, each contributing uniquely to the joke’s overall impact.
Pun usage demonstrated selective associations, showing moderate correlations with sound-based features including rhyme (Pearson: 0.29, Spearman: 0.26) and homonyms (Pearson: 0.28, Spearman: 0.20). However, correlations with other mechanisms were weaker (Pearson(r) <0.25, Spearman (r) <0.10). This pattern suggests that while puns integrate sound patterns and multiple meanings, they operate through relatively distinct linguistic pathways in humour generation.
Mechanism Correlation Conclusion
The comparison between Pearson and Spearman correlations reveals that humour mechanisms interact through multiple parallel pathways rather than a single unified system. While some mechanisms show strong linear relationships in extreme cases (as evidenced by high Pearson correlations), the lower Spearman correlations suggest these relationships are not typical of most humorous content. We observe the following:
  • Outlier-driven relationships: Particularly between semantic conflicts and rhyme
  • Stable moderate correlations: Especially in homonym-based features
  • Independent mechanisms: Such as alliteration and self-reference
  • Context-dependent interactions: Varying strength of relationships based on humour style and content
This analysis suggests that effective humour can be constructed through various combinations of these mechanisms, with no single pattern dominating across all instances.